# BOOST: Bayesian Optimization with Optimal Kernel and Acquisition Function Selection Technique

### v1

@@ Abstract
The performance of Bayesian optimization (BO), a highly sample-efficient method for expensive black-box problems, is critically governed by the selection of its hyperparameters, including the kernel and acquisition functions. This presents a significant practical challenge: an inappropriate combination of these can lead to poor performance and wasted evaluations. While individual improvements to kernel functions (e.g., tree-based kernels, deep kernel learning) and acquisition functions (e.g., multi-step lookahead, tree-based planning) have been actively explored, the joint and autonomous selection of the best pair of these fundamental hyperparameters has been largely overlooked. This forced practitioners to rely on heuristics or costly manual training. In this work, we propose a simple yet effective framework, BOOST (Bayesian Optimization with Optimal Kernel and Acquisition Function Selection Technique), that automates this selection. BOOST utilizes a lightweight, offline evaluation stage to predict the performance of various kernel-acquisition function pairs and identify the most suitable configuration before committing to the expensive evaluation process. BOOST partitions data-in-hand into two subsets: a reference subset and a query subset, and it prepares all possible kernel-acquisition pairs from the user’s chosen candidates. For each configuration, BOOST conducts internal BO runs using the reference subset, evaluating how effectively each pair guides the search toward the optimum in the unknown query subset, thereby identifying the configuration with the best retrospective performance for future optimization. Experiments on both synthetic benchmark functions and real-world hyperparameter optimization tasks demonstrate that BOOST consistently outperforms standard BO approaches with fixed hyperparameters, highlighting its effectiveness and robustness in diverse problem landscapes.
Introduction
@@
Black-box optimization is a common challenge in scientific and engineering problems, especially when each function evaluation is time-consuming or costly (Jones et al. 1998). In this context, a black box refers to situations where the objective function cannot be modeled and only input-output relationships are observable, with no access to gradients or knowledge of the underlying system. These challenges necessitate efficient methods that minimize the number of evaluations to approach optimal solutions.
@@
The fundamental challenge in black-box optimization lies in managing uncertainty about the objective function’s behavior in unexplored regions. The high cost of function evaluations means that only a limited number of observations can be obtained, requiring optimization strategies to make informed decisions under uncertainty with sparse data. Bayesian Optimization (BO) addresses this challenge by employing a principled probabilistic approach that combines a surrogate model—most commonly a Gaussian Process (GP) (Williams and Rasmussen 2006)—with an acquisition function to guide the search process efficiently (Frazier 2018; Shahriari et al. 2015). GP provides uncertainty estimates of the objective function, while the acquisition function leverages these estimates to determine the next promising point to evaluate. This combination allows BO to systematically balance exploration and exploitation, leading to data-efficient optimization (Snoek et al. 2012).
@@
Given these advantages, BO is broadly used for costly experiments, including material synthesis (Pruksawan et al. 2019; Shields et al. 2021), optimizing process variables (Greenhill et al. 2020; Park et al. 2024a, b), and machine learning hyperparameter optimizations (Chen et al. 2018; Falkner et al. 2018; Feurer et al. 2015; Kandasamy et al. 2018; Shen et al. 2023; Snoek et al. 2012; White et al. 2021).
@@
Despite its widespread application, the performance of BO critically depends on the proper choice of two hyperparameters: the choice of kernel function used in the GP and the choice of acquisition function guiding the subsequent sample point (Ginsbourger et al. 2008; Hoffman et al. 2011; Roman et al. 2019; Snoek et al. 2012; Vasconcelos et al. 2019). Nevertheless, these hyperparameters are typically fixed throughout the optimization process, generally chosen arbitrarily or based on prior domain knowledge (Genton 2001; Wilson and Adams 2013), with little attention paid to their selection or adaptation (Roman et al. 2019). This oversight neglects an important insight: the optimal combination of kernel and acquisition function can vary significantly depending on the structure and complexity of the objective function. However, the black-box nature of these problems creates a fundamental paradox: practitioners must select hyperparameters that match the function’s characteristics without knowing those characteristics beforehand.
@@
Some existing approaches have attempted to address this gap by adaptive selection of kernel functions (Ginsbourger et al. 2008; Roman et al. 2019) or acquisition functions (Hoffman et al. 2011; Vasconcelos et al. 2019). However, these methods typically rely on local information, such as the outcome of a single query point or uncertainty estimates from the surrogate model, which may not always reflect the true structure of the objective function, particularly in noisy or complex optimization landscapes. Another active research direction is individual development of kernel functions (Wilson et al. 2016; Boyne et al. 2025) or acquisition functions (Wu and Frazier 2019; Lam et al. 2016; Song et al. 2022; Cheon et al. 2024). These methods, however, focus on improving either the kernel or the acquisition function while leaving the other fixed, without considering their joint selection.
@@
To address these challenges, we propose BOOST (Bayesian Optimization with Optimal Kernel and Acquisition Function Selection Technique), a simple yet effective method that identifies the most suitable kernel–acquisition pair based entirely on the previously evaluated points. By performing offline evaluations using the data-in-hand, BOOST avoids dependence on uncertain model estimates and local improvements.
@@
For each BO iteration, BOOST selects the suitable kernel-acquisition function pair through a three-stage process: data partitioning, performance evaluation, and recommendation. We first partition the data-in-hand into two subsets: one used as the reference set to construct a GP for the internal BO process, and the other treated as a query set which serves as unexplored regions of internal BO where new query points are generated.
For each candidate kernel-acquisition pair predefined by users, BOOST fits a GP to the reference subset and selects the next query point in the query subset. This internal BO is repeated until it reaches the target value in the query subset. The number of iterations required by each pair to reach the target is recorded, and the pair with minimum iteration number is selected for the actual BO iteration.
@@
Preliminaries
Assume we aim to find the global maximum,
$x^{*}=\arg\max_{x\in\mathcal{X}}f(x),$
@@
$k_{\mathrm{RBF}}(r)=\exp\left(-\frac{r^{2}}{2l^{2}}\right)$ (6)
$k_{\mathrm{RQ}}(r)=\left(1+\frac{r^{2}}{2\alpha l^{2}}\right)^{-\alpha}$ (7)
@@
The Radial Basis Function (RBF) kernel (Lowe and Broomhead 1988) assumes an infinitely smooth function. In contrast, the Matérn kernel (Matérn 1960; Williams and Rasmussen 2006) offers flexibility in modeling less smooth functions, with $\nu=$ 3/2 and 5/2 being widely used for real-world tasks. The Rational Quadratic (RQ) kernel (Williams and Rasmussen 2006) is capable of modeling functions with multiple characteristic length scales, with $\alpha=2$ being widely used. Throughout this paper, we fix the parameter $\alpha$ of the RQ kernel to 2. Given this variability, the choice of kernel can have a substantial impact on BO performance depending on the structure of the objective function.
Acquisition function
@@
Among various acquisition functions, Expected Improvement (EI) (Jones et al. 1998) is one of the most widely used. It computes the expected improvement over the current best observation, thereby naturally balancing exploration (through predictive uncertainty) and exploitation (through mean estimate). The following are widely used acquisition functions.
$\alpha_{\mathrm{EI}}(x)=\mathbf{E}\left[\max(f(x)-f^{+},0)\right]$ (8)
$\alpha_{\mathrm{PI}}(x)=\mathbf{P}(f(x)>f^{+})$ (9)
$\alpha_{\mathrm{UCB}}(x)=\mu(x)+\beta\sigma(x)$ (10)
$\alpha_{\mathrm{PM}}(x)=\mu(x)$ (11)
where $f^{+}$ denotes the best observed function value so far.
@@
Different acquisition functions correspond to different search strategies. Probability of Improvement (PI) (Kushner 1964) focuses solely on the probability of achieving better outcomes than the current best, often leading to overly greedy behavior if uncertainty is underestimated. Upper Confidence Bound (UCB) (Srinivas et al. 2010) adds a weighted uncertainty term to the predicted mean, enabling more principled exploration. The parameter $\beta$ controls the exploration–exploitation trade-off: a low $\beta$ (e.g., 0.1) encourages exploitation behavior, often more than PI, while a high $\beta$ (e.g., 5) encourages aggressive exploration, surpassing even EI in exploratory behavior (Papenmeier et al. 2025). Since no standard value exists for $\beta$, we adopt $\beta=0.1$ with reference to the values used in some recent studies (Dogan and Prestwich 2022; Tian et al. 2024). In contrast, Posterior Mean (PM) chooses the next point solely based on the GP’s predicted mean value. While this may work well when the model is highly confident, it completely ignores uncertainty and thus represents the most greedy strategy among common acquisition functions.
Related Works
Adaptive Kernel and Acquisition Selection in BO
@@
Various approaches have been proposed to address the challenge of selecting appropriate kernel and acquisition functions in Bayesian Optimization (BO), recognizing their strong influence on performance. GP-Hedge (Hoffman et al. 2011) addresses the acquisition selection problem by treating it as a multi-armed bandit, adaptively adjusting the probability distribution over acquisition functions based on the surrogate model’s predicted value at each candidate point they propose. However, this approach heavily relies on uncertain, model-driven improvements from single points, leading to potential instability, especially in the early stages of optimization. No-PASt-BO (Vasconcelos et al. 2019) mitigates this by incorporating historical data with a memory factor, yet still suffers from reliance on uncertain predictive values and does not address kernel adaptation.
@@
Similarly, adaptive kernel selection approaches have been explored. Ginsbourger et al. (2008) proposed using discrete mixtures of kernels within Gaussian Process-based optimization (originally referred to as Kriging), where the surrogate model is constructed from a weighted combination of multiple kernels. Roman et al. (2019), on the other hand, developed a strategy to select the best-performing kernel among multiple candidates by evaluating them in parallel. While these approaches reduce kernel misspecification risks, they both rely on local surrogate information for decision-making and neglect the influence of acquisition function choice, which can limit their overall robustness and adaptability in complex optimization settings.
@@
In contrast to these approaches that focus on individual components, some recent work has recognized the need for joint selection. The fundamental challenge in black-box optimization is that practitioners must choose appropriate kernel and acquisition functions that match the function’s characteristics without knowing those characteristics beforehand, making simultaneous optimization of both components particularly difficult. To address this issue, Xue et al. (2016) proposed a data-based strategy that utilizes the data-in-hand to jointly select the most appropriate surrogate model and acquisition function for material design tasks. While this approach represents a significant step toward breaking the conventional paradox of selecting hyperparameters without prior knowledge—by transforming available data into actionable insights—their method selects hyperparameters only at the beginning and does not update them during subsequent iterations. This static approach fails to leverage the additional information that becomes available as the optimization progresses, limiting its adaptability to the evolving understanding of the objective function.
@@ Methods
Overall Process of BOOST
@@
BOOST performs offline evaluation using data-in-hand to identify the most suitable kernel-acquisition pair from candidates predefined by practitioners. BOOST operates in three main steps:
@@
    Data Partitioning: The data-in-hand, $D_{n}$, is partitioned into two disjoint subsets: a reference subset, $r_{n}$, to construct the Gaussian Process (GP) for the internal BO process, and a query set, and $q_{n}$ which serves as unexplored regions of internel BO where new query points are generated.
@@
    Performance Evaluation: The optimization performance of each candidate kernel–acquisition pair is evaluated based on how quickly it discovers the target value $y_{target}$ within $q_{n}$, starting from $r_{n}$.
@@
The overview of BOOST is shown in Figure 1 and Algorithm 1.
Algorithm 1 Overall algorithm for BOOST
Output: Best Kernel-Acquitision function pair ($k^{*}$, $\alpha^{*}$)
[…]
5:for all $(k_{i},\alpha_{j})$ in $\mathcal{K}\times\mathcal{A}$ do
6:  $R_{0}\leftarrow r_{n}=\{(x_{i},f(x_{i}))\}_{i=1}^{n_{init}}$
7:  $Q_{0}\leftarrow\{x\mid(x,f(x))\in q_{n}\}$
8:  $t\leftarrow 0$
9:  while $t<T_{\max}$ do
10:   Train a Gaussian Process using $k_{i}$ over $R_{t}$
11:   $x_{t+1}\leftarrow\arg\max_{x\in Q_{t}}\alpha_{j}(x)$
12:   $R_{t+1}\leftarrow R_{t}\cup\{(x_{t+1},f(x_{t+1}))\}$
13:   $Q_{t+1}\leftarrow Q_{t}\setminus\{x_{t+1}\}$
14:   $t\leftarrow t+1$
15:   if $f(x_{t+1})\leq y_{target}$ then
16:    break
17:   end if
18:  end while
19:  $t_{k_{i},\alpha_{j}}\leftarrow t$
20:end for
Data Partitioning
@@
To ensure that $r_{n}$ consists of informative and diverse samples while excluding points already better than the target value (or stopping criteria), $y_{target}$, we apply K-means clustering on $D_{n}\setminus\{(x,f(x))\mid f(x)\leq y_{target}\}$ to select representative samples that provide diverse coverage of the promising regions. From each cluster, the data point closest to the cluster center is selected, providing both representativeness and coverage of promising regions. In this study, we define the target value $y_{target}$ for stopping as the top 5th percentile of function values in $D_{n}$, while the size of $r_{n}$ is determined as $\lfloor|D_{n}|/3\rfloor$ with the constraint of $3\leq|r_{n}|\leq 20$ to balance the size of $r_{n}$ and $q_{n}$, where $q_{n}=D_{n}\setminus r_{n}$. The performance impact of different stopping criteria and data partitioning methods is analyzed in the Appendix.

### v2

@@ Abstract
The performance of Bayesian optimization (BO), a highly sample-efficient method for expensive black-box problems, is critically governed by the selection of its hyperparameters, including the kernel and acquisition functions. This presents a significant practical challenge: an inappropriate combination of these can lead to poor performance and wasted evaluations. While individual improvements to kernel functions (e.g., tree-based kernels, deep kernel learning) and acquisition functions (e.g., multi-step lookahead, tree-based planning) have been actively explored, the joint and autonomous selection of the best pair of these fundamental hyperparameters has been largely overlooked. This forced practitioners to rely on heuristics or costly manual training. In this work, we propose a highly novel framework, BOOST (Bayesian Optimization with Optimal Kernel and Acquisition Function Selection Technique), that automates this selection. BOOST utilizes a lightweight, offline evaluation stage to predict the performance of various kernel-acquisition function pairs and identify the most promising pair before committing to the expensive evaluation process. Using K-means clustering, BOOST first chooses initial subsets among previously observed points (data-in-hand), and it prepares all possible kernel-acquisition pairs from the user’s chosen candidates. For each pair, BOOST conducts internal BO runs starting with initial subset, evaluating how many iterations has been required to find the target value within the remaining data, thereby identifying the kernel-acquisition pair with the best retrospective performance for future optimization. Experiments on both synthetic benchmark functions and real-world hyperparameter optimization tasks demonstrate that BOOST consistently outperforms standard BO approaches with fixed hyperparameters as well as state-of-the-art adaptive methods, highlighting its effectiveness and robustness in diverse problem landscapes.
1 Introduction
@@
Black-box optimization is a common challenge in scientific and engineering problems, especially when each function evaluation is time-consuming or costly (Jones et al.,, 1998). In this context, a black box refers to situations where the objective function cannot be modeled and only input-output relationships are observable, with no access to gradients or knowledge of the underlying system. These challenges necessitate efficient methods that minimize the number of evaluations to approach optimal solutions.
@@
The fundamental challenge in black-box optimization lies in managing uncertainty about the objective function’s behavior in unexplored regions. The high cost of function evaluations means that only a limited number of observations can be obtained, requiring optimization strategies to make informed decisions under uncertainty with sparse data. Bayesian Optimization (BO) addresses this challenge by employing a principled probabilistic approach that combines a surrogate model—most commonly a Gaussian Process (GP) (Williams and Rasmussen,, 2006)—with an acquisition function to guide the search process efficiently (Shahriari et al.,, 2015; Frazier,, 2018). GP provides uncertainty estimates of the objective function, while the acquisition function leverages these estimates to determine the next promising point to evaluate. This combination allows BO to systematically balance exploration and exploitation, leading to data-efficient optimization (Snoek et al.,, 2012).
@@
Given these advantages, BO is broadly used for costly experiments, including material synthesis (Pruksawan et al.,, 2019; Shields et al.,, 2021), optimizing process variables (Greenhill et al.,, 2020; Park et al., 2024a, ; Park et al., 2024b, ), and machine learning hyperparameter optimizations (Snoek et al.,, 2012; Klein et al.,, 2017; Chen et al.,, 2018; Falkner et al.,, 2018; Kandasamy et al.,, 2018).
@@
Despite its widespread application, the performance of BO critically depends on the proper choice of two hyperparameters: the choice of kernel function used in the GP and the choice of acquisition function guiding the subsequent sample point (Ginsbourger et al.,, 2008; Hoffman et al.,, 2011; Snoek et al.,, 2012; Roman et al.,, 2019; Vasconcelos et al.,, 2019). Nevertheless, these hyperparameters are typically fixed throughout the optimization process, generally chosen arbitrarily or based on prior domain knowledge (Genton,, 2001; Wilson and Adams,, 2013), with little attention paid to their selection or adaptation (Roman et al.,, 2019). This oversight neglects an important insight: the optimal combination of kernel and acquisition function can vary significantly depending on the structure and complexity of the objective function. However, the black-box nature of these problems creates a fundamental paradox: practitioners must select hyperparameters that match the function’s characteristics without knowing those characteristics beforehand.
@@
Some existing approaches have attempted to address this gap by adaptive selection of kernel functions (Ginsbourger et al.,, 2008; Roman et al.,, 2019) or acquisition functions (Hoffman et al.,, 2011; Vasconcelos et al.,, 2019). However, these methods typically rely on local information, such as the outcome of a single query point or uncertainty estimates from the surrogate model, which may not always reflect the true structure of the objective function, particularly in noisy or complex optimization landscapes. Another active research direction is individual development of kernel functions (Wilson et al.,, 2016; Eriksson and Jankowiak,, 2021; Jimenez and Katzfuss,, 2023; Boyne et al.,, 2025) or acquisition functions (Lam et al.,, 2016; Wu and Frazier,, 2019; Song et al.,, 2022; Astudillo et al.,, 2023; Cheon et al.,, 2024). These methods, however, focus on improving either the kernel or the acquisition function while leaving the other fixed, without considering their joint selection.
@@
To address these challenges, we propose BOOST (Bayesian Optimization with Optimal Kernel and Acquisition Function Selection Technique), a highly novel framework that identifies the most promising kernel–acquisition pair based entirely on previously evaluated points. By performing offline evaluations using previously observed points (data-in-hand), BOOST avoids dependence on uncertain model estimates and local improvements.
@@
For each BO iteration, BOOST selects the most promising kernel-acquisition function pair through a three-stage process: data partitioning, performance evaluation, and recommendation. We first partition the data-in-hand into two subsets using K-means clustering. The data points closest to each cluster center serve as the initial subset for constructing a GP (reference set), and the remainder serves as unexplored regions where new query points are generated (query set).
In the performance evaluation step, for each candidate kernel-acquisition pair, BOOST conducts internal BO runs starting with the reference set, evaluating how many iterations have been required to find the target value within the query set. The pair requiring minimum iterations is selected for the actual BO iteration in the recommendation step.
@@
2 Preliminaries
Assume we aim to find the global minimum,
$x^{*}=\arg\min_{x\in\mathcal{X}}f(x),$
@@ 2.1 Gaussian Process (GP)
$k_{\mathrm{RBF}}(r)=\exp\left(-\frac{r^{2}}{2\ell^{2}}\right)$ (6)
$k_{\mathrm{RQ}}(r)=\left(1+\frac{r^{2}}{2\alpha\ell^{2}}\right)^{-\alpha}$ (7)
@@
The Radial Basis Function (RBF) kernel (Lowe and Broomhead,, 1988) assumes an infinitely smooth function. In contrast, the Matérn kernel (Matérn,, 1960; Williams and Rasmussen,, 2006) offers flexibility in modeling less smooth functions, with $\nu=$ 3/2 and 5/2 being widely used for real-world tasks. The Rational Quadratic (RQ) kernel (Williams and Rasmussen,, 2006) is capable of modeling functions with multiple characteristic length scales, with $\alpha=2$ being widely used. Throughout this paper, we fix the parameter $\alpha$ of the RQ kernel to 2. Given this variability, the choice of kernel can have a substantial impact on BO performance depending on the structure of the objective function.
2.2 Acquisition function
@@
Among various acquisition functions, Expected Improvement (EI) (Jones et al.,, 1998) is one of the most widely used. It computes the expected improvement over the current best observation, thereby naturally balancing exploration (through predictive uncertainty) and exploitation (through mean estimate). The following are widely used acquisition functions.
$\alpha_{\mathrm{EI}}(x)=\mathbf{E}\!\left[\max\!\big(f^{-}-f(x),\,0\big)\right],$ (8)
$\alpha_{\mathrm{PI}}(x)=\mathbf{P}\big(f(x)<f^{-}\big),$ (9)
$\alpha_{\mathrm{LCB}}(x)=-(\mu(x)-\beta\sigma(x)),$ (10)
$\alpha_{\mathrm{PM}}(x)=-\mu(x),$ (11)
where $f^{-}$ denotes the lowest observed value so far. Note that maximizing Eq. (10) results in $x_{n+1}=\arg\min_{x}\mu(x)-\beta\sigma(x)$, and maximizing Eq. (11) results in $x_{n+1}=\arg\min_{x}\mu(x)$.
@@
Different acquisition functions correspond to different search strategies. Probability of Improvement (PI) (Kushner,, 1964) focuses solely on the probability of achieving better outcomes than the current best, often leading to overly greedy behavior if uncertainty is underestimated. Lower Confidence Bound (LCB), which is the minimization version of Upper Confidence Bound (Srinivas et al.,, 2010), adds a weighted uncertainty term to the predicted mean, enabling more principled exploration. The parameter $\beta$ controls the exploration–exploitation trade-off: a low $\beta$ (e.g., 0.1) encourages exploitation behavior, often more than PI, while a high $\beta$ (e.g., 5) encourages aggressive exploration, surpassing even EI in exploratory behavior (Papenmeier et al.,, 2025). Since no standard value exists for $\beta$, we adopt $\beta=0.1$ with reference to the values used in some recent studies (Dogan and Prestwich,, 2022; Tian et al.,, 2024). In contrast, Posterior Mean (PM) chooses the next point solely based on the GP’s predicted mean value. While this may work well when the model is highly confident, it completely ignores uncertainty and thus represents the most greedy strategy among common acquisition functions.
3 Related Works
@@
Various approaches have been proposed to address the challenge of selecting appropriate kernel and acquisition functions in Bayesian Optimization (BO), recognizing their strong influence on performance. GP-Hedge (Hoffman et al.,, 2011) addresses the acquisition selection problem by treating it as a multi-armed bandit, adaptively adjusting the probability distribution over acquisition functions based on the surrogate model’s predicted value at each candidate point they propose. However, this approach heavily relies on uncertain, model-driven improvements from single points, leading to potential instability, especially in the early stages of optimization. No-PASt-BO (Vasconcelos et al.,, 2019) mitigates this by incorporating historical data with a memory factor, yet still suffers from reliance on uncertain predictive values and does not address kernel adaptation.
@@
Similarly, adaptive kernel selection approaches have been explored. Ginsbourger et al., (2008) proposed using discrete mixtures of kernels within Gaussian Process-based optimization (originally referred to as Kriging), where the surrogate model is constructed from a weighted combination of multiple kernels. Roman et al., (2019), on the other hand, developed a strategy to select the best-performing kernel among multiple candidates by evaluating them in parallel. While these approaches reduce kernel misspecification risks, they both rely on local surrogate information for decision-making and neglect the influence of acquisition function choice, which can limit their overall robustness and adaptability in complex optimization settings.
@@
In contrast to these approaches that focus on individual components, some recent work has recognized the need for joint selection. The fundamental challenge in black-box optimization is that practitioners must choose appropriate kernel and acquisition functions that match the function’s characteristics without knowing those characteristics beforehand, making simultaneous optimization of both components particularly difficult. To address this issue, Xue et al., (2016) proposed a data-based strategy that utilizes the data-in-hand to jointly select the most appropriate surrogate model and acquisition function for material design tasks. While this approach represents a significant step toward breaking the conventional paradox of selecting hyperparameters without prior knowledge—by transforming available data into actionable insights—their method selects hyperparameters only at the beginning and does not update them during subsequent iterations. This static approach fails to leverage the additional information that becomes available as the optimization progresses, limiting its adaptability to the evolving understanding of the objective function.
@@ 4 Methods
4.1 Overall Process of BOOST
@@
BOOST performs offline evaluation using previously observed points (data-in-hand) to identify the most promising kernel-acquisition pair from candidates predefined by practitioners. BOOST operates in three main steps:
@@
    Data Partitioning: The data-in-hand, $D_{n}$, is partitioned into two disjoint subsets using K-means clustering. The data points closest to each cluster center serve as the initial subset for constructing a GP (reference set, $r_{n}$), and the remainder serves as unexplored regions where new query points are generated (query set, $q_{n}$).
@@
    Performance Evaluation: For each candidate in kernel-acquisition pairs, BOOST conducts internal BO runs starting with $r_{n}$. The optimization performance is evaluated based on how many iterations have been required to find the target value, $y_{target}$, within $q_{n}$.
@@
The overview of BOOST is shown in Figure 1 and Algorithm 1 in Appendix C.
4.2 Data Partitioning
@@
To ensure that $r_{n}$ consists of informative and diverse samples while excluding points already better than the target value (or stopping criteria), $y_{target}$, we apply K-means clustering on $D_{n}\setminus\{(x,f(x))\mid f(x)\leq y_{target}\}$ to select representative samples that provide diverse coverage of the promising regions. From each cluster, the data point closest to the cluster center is selected, providing both representativeness and coverage of promising regions. In this study, we define the target value $y_{target}$ for stopping as the top 5th percentile of function values in $D_{n}$, while the size of $r_{n}$ is determined as $\lfloor|D_{n}|/3\rfloor$ with the constraint of $3\leq|r_{n}|\leq 20$ to balance the size of $r_{n}$ and $q_{n}$, where $q_{n}=D_{n}\setminus r_{n}$. The performance impact of different stopping criteria and data partitioning methods is analyzed in Appendix F.